Subject: 3.5-m Users Committee minutes 6/11/07

From: Bruce Gillespie

Submitted: Wed, 13 Jun 2007 13:55:26 -0600

Message number: 1032 (previous: 1031, next: 1033 up: Index)

                         APO 3.5-m Users Committee Phonecon, 6/11/07

Attending:  Suzanne Hawley, Bruce Gillespie, Mark Klaene, Michael  
Strauss, Jon Holtzman, John Bally, Al Harper

Absent: Scott Anderson, Jon Fulbright, Remy Indebetouw

Minutes taken by Bruce Gillespie


User feedback, comments from institutional representatives:

Princeton (Strauss) - Michael said everything is going fine with the  
PU users.

JHU (Fulbright) - no report.

Chicago (Harper) - Nothing to report from the users. Al was  
interested in the status/plans for the upcoming NIC-FPS upgrades.   
John said that the engineering detector is now at Rockwell for  
modifications to enable the 16-channel readouts (the science-grade  
detector will get the same treatment during the shutdown), and that  
they are waiting for the test dewar to be delivered to CU.  The plan  
is to take NIC-FPS to Boulder in the third week of July, and if  
things go well with the detector upgrades and Fabry-Perot work, the  
instrument may be returned to APO near the third week in September.   
The return date should be better known in early August, and users  
should be aware that NIC-FPS availability in the 3rd quarter is  
somewhat uncertain for the moment.

Washington (Suzanne) - Suzanne reported that UW observers had no  
problems other than weather-related.  Lucianne Walkowicz did get some  
decent observing weather and was able to finish her thesis data  

NMSU (Holtzman) - Jon had nothing to report, other than weather woes.

Virginia (Indebetouw) - no report.

Colorado (Bally) - John said that things were going OK.


Discussion of telescope/instruments report - Klaene:

The report is given below; Mark walked through the highlights.  All  
three of the old, refurbished drive boxes were installed on the  
telescope with minimal lost observing time, and the telescope  
pointing/tracking performance has shown improvement.  The new boxes  
that were removed had been in service for about a year and had begun  
to show signs of wear after 9 months.  They will be refurbished in  
reinstalled next year.  Everything is on schedule for the summer  
shutdown (dates below); the main item on the shutdown is the  
realuminization of the primary mirror at KPNO.

For the new TUI/SPIcam software, it is being test-driven by users and  
we are collecting feedback.  We're still working on scripting and sub- 
framing issues.  We expect that TUI/SPIcam will be the default for  
users in Q3.  In order to use SPIcam with TUI, users need to install  
a new version of TUI.  This will be announced with the general  
availability information after the user tests.

Mark added that since the report was written, the power supply for  
the echelle quartz flat-field lamp had failed.  We expect a  
replacement this week, and in the meantime recent archive flats have  
been furnished to echelle users.

On the CIF projects, Mark mentioned the redesign of the tertiary  
mirror supports, which will improve stability and maintainability.  A  
tip-tilt capability is under discussion for the tertiary rebuild, and  
we are looking for science- and experience-based input and advice on  
whether to include hooks in the design for tip-tilt actuators.    
Hooks in the top-end rebuild have been left for tip-tilt with the  
secondary mirror, but it is unclear whether doing tip-tilt with the  
secondary or tertiary mirrors is preferable, if we do it at all.   
Anyone who has experience and/or interest in tip-tilt and its  
implementation on the 3.5-m and would like to contribute to this  
discussion should contact Suzanne and Bruce right away, as we will be  
deciding on the tertiary design parameters this summer.  Also, a  
project has begun to design a direct-drive system for the alt and az  
motions on the telescope to improve performance and maintainability.   
Lastly, a mirror-cell port instrument rotator is being considered,  
first for Agile and then as a general-purpose rotator for small  


    3.5-m Telescope, Instruments, and CIF Projects Highlights,  
5/10/07 through 6/8/07

                                  Mark Klaene

0) Overview
Spring weather continues to be intermixed with cloud cover and light  
precipitation.  Fire danger is unlikely to reach extreme levels this  
year, but is increasing.  Some fir trees near the observatory  
attacked by defoliators are showing signs of foliage regeneration.   
We are working with the USFS on possible treatments.

1) Telescope
No significant problems this period.  Engineering time in May was  
used to swap out azimuth and altitude drive boxes and service the  
main azimuth bearing.  Tracking performance is very good coming out  
of this shutdown.

The summer shutdown will be from August 20 through September 16.  The  
schedule driver is the re-aluminization of the primary mirror.  We  
have a limited number of days on the sky for recovery; hopefully the  
weather will cooperate.

2) Instruments

Agile is preparing for its next commissioning run in July.  The  
camera recently went back to Princeton Instruments to get the fiber- 
optic interface to work.

The modified grating holder (version 2) for the DIS gratings was  
tested and found to be safer to install and still be repeatable.  The  
new holder has mid-size kinematic bushings that will be retrofitted  
into all grating holders this summer.

The replacement DIS red high-resolution grating has been received and  
is at JHU for testing.  We expect delivery to the observatory this week.

The flip mirror motor on the Echelle was replaced.

Incorporation of SPICam into TUI (TUI 1.4.a5) has been released for  
initial user testing.  Some user scripts are still in work.

The NA2 Guider camera had an intermittent network connection failure  
develop. The cause was traced to a broken cable retainer clip which  
has since been repaired.

3) Engineering and CIF projects


Final quotes have been obtained for the rotator upgrade and we are  
issuing purchase orders.

Bench setup has been built up and we will begin testing the new axis  
controllers this week.


Plans are underway for two major telescope projects:

The first involves an upgrade to the tertiary mounting.  This will  
stiffen the tertiary mounting to improve pointing and improve  
reliability.  Discussion is ongoing as to the need for tip/tilt to be  

Secondly, we are pursuing a design for putting direct drives on the  
azimuth and altitude axes.  This can be implemented only after the  
new axis controllers are ready for the axes.  Direct drives would  
improve tracking performance and stability and reduce maintenance and  
increase reliability.  Some improvement in slew speeds could also be  
realized.  We are going to get some experience with a direct drive  
system in the 2007 rotator upgrade project, which will be useful in  
the design and implementation of direct drives for the telescope alt  
and az axes.

4) Miscellaneous

Jim Davenport will be at APO starting this week as an extra Observing  
Specialist to help fill in during summer staff absences.


New SPIcam/TUI interface - Group:

[see discussion above]


Interest in a new imaging camera - Hawley:

John Bally said he'd be very interested in a wider-field, visible  
imager with a new detector.  Suzanne mentioned that there is  
discussion of a new NSF program for medium-sized telescopes that  
could enable either a new camera for us, or access to comparable  
telescopes with wide-field imagers (e.g., the ODI on WIYN).  The  
question for our community is whether or not there is sufficient  
interest to go after a new imager or access to one on another  
telescope.  Bruce said that there still is an open question of if  
there is an existing wide-field corrector for the 3.5-m, one that was  
purportedly built by Chris Stubbs but seemingly lost in storage  
somewhere.  John Bally said that when he previously looked into wide- 
field imaging, he found that some kind of field correction would be  
needed if we were to image fields 10-arcminutes or larger.

With respect to the NSF mid-sized telescope system program mentioned  
above, shortly after the users committee meeting, the solicitation  
for input to the ReSTAR committee that is charged by NSF with  
documenting the
science case for 1- to 5-meter class telescopes in the next decade  
was released.  Suzanne forwarded this solicitation to apo35-general.   
Users should go to the ReSTAR website at: 
restar and fill in the web-based form to provide input about how the  
3.5m telescope (and other 1- to 5-meter telescopes) are useful for  
their research.


Actions from earlier meetings - group

[open from previous months]:


[new actions from last month]:



Next meeting:

The next Users Committee phonecon will be on Monday, 16 July, at 8:30  
AM Pacific Time.  Agenda and other materials will be sent to the  
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1032 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to